OpenStudy (anonymous):

Use the formula for future value, A = P(1 + rt), to find the missing quantity. A = $8,680; P = $7,000; r = 8%

OpenStudy (kirbykirby):

\[8680=7000(1+0.08t)\\ \frac{8680}{7000}=1+0.08t\\ 1.24-1=0.08t\\ \frac{0.24}{0.08}=t\]

OpenStudy (anonymous):

thank you

OpenStudy (kirbykirby):

=]